OpenStudy (anonymous):

Find the slope-intercept equation of the line that has the given characteristics: Slope 1.7 and y-intercept (0, -6). y = ?

OpenStudy (aroub):

y=mx+b -6=1.7(0)+b -6=b y=1.7x

jimthompson5910 (jim_thompson5910):

aroub, your last step (ie your final answer) should be y=1.7x-6

OpenStudy (aroub):

yes!! sorryyy

jimthompson5910 (jim_thompson5910):

np silly typo happens to me all the time